# On stable assignments generated by choice functions of mixed type

### v1

@@ 3 Finding an initial stable assignment
In order to construct the stable matching $x^{\rm min}$ (which is the best for the part $F$ and the worst for the part $W$) in our mixed problem SMP, one can use two approaches. One of them follows a method applied in [1, Sec. 3.1] to show the existence of a stable assignment in a general case considered there. The other one generalizes ideas of the method of finding the initial stable allocation elaborated in [4].
We prefer to describe the first method, which is shorter and looks more enlightening. It iteratively constructs triples of functions $(b^{i},x^{i},y^{i})$, $i=0,1,\ldots,i,\ldots$  .
@@
At the first stage of the iteration, $b^{i}$ is transformed into $x^{i}$, by putting $x^{i}_{f}:=C_{f}(b^{i}_{f})$ for each $f\in F$. (Then $x^{i}$ satisfies the quotas $q(f)$ for all “firms” $f\in F$, however, it can exceed quotas $q(w)$ for some “workers” $w\in W$.)
At the second stage, the obtained $x^{i}$ is transformed into $y^{i}$, by putting $y^{i}_{w}:=C_{w}(x^{i}_{w})$ for each $w\in W$. (This $y^{i}$ already satisfies the quotas $q(v)$ for all $v\in V$, but it need not be stable.)
The obtained $x^{i}$ and $y^{i}$ are then used to generate the upper bounds $b^{i+1}$ for the next iteration, by putting
@@
The process terminates if at some iteration, $i$-th say, we obtain the equality $y^{i}=x^{i}$ (which implies $b^{i+1}=b^{i}$, by (3.1)); then $x^{i}$ is just the required assignment (see the proposition below). In reality, it may happen that the sequence of iterations is infinite. Nevertheless, one can see (by arguing as in [1]) that the process converges. Indeed, the construction shows that for each $i$,
@@
This result was proved by Alkan and Gale for a general case considered in [1]; see Theorem 1 (for the stability) and Theorem 2 (for the optimality) in that paper.
Fortunately, for our mixed model the above method can be modified, by aggregating some sets of consecutive iterations, so as to obtain an efficient algorithm to find the assignment $x^{\rm min}$; its running time is polynomial in $|E|$. We describe this modified method for SMP in the Appendix.
@@ 4 Rotations
Throughout this section we consider a fixed stable assignment $x\in{\cal S}$ for our mixed problem SMP. Our aim is to characterize the set of stable assignments $x^{\prime}$ that are close to $x$ and satisfy $x\succ_{F}x^{\prime}$. This will be done by use of a technique of rotations, which generalizes that developed for the stable allocation problem in [4]. First of all we specify some notation needed in what follows.
$\bullet$ We denote by $F^{=}$ and $W^{=}$ the sets of fully filled vertices in the parts $F$ and $W$, respectively, that is, the vertices $f\in F$ and $w\in W$ such that $|x_{f}|=q(v)$ and $|x_{w}|=q(w)$. (By ((2.7))(b), these sets do not depend on the choice of $x\in{\cal S}$. Furthermore, in view of ((2.8)), any transformation $x\rightsquigarrow x^{\prime}\in{\cal S}$ should involve only edges connecting $F^{=}$ and $W^{=}$.)
@@
$\bullet$ We say that an edge $e\in E$ is saturated by $x$ if $x(e)=b(e)$; the set of such edges is denoted as $U(x)$, and its restriction to $E_{v}$ ($v\in V$) as $U_{v}(x)$.
$\bullet$ Usually we will abbreviate $\succeq_{F}$ to $\succeq$.
@@
    for each non-saturated edge $e=fw\in E-U(x)$, at least one vertex $v$ among $f$ and $w$ is such that: $v$ is fully filled and $e$ belongs to $H_{v}(x)\cup E^{\rm end}_{v}(x)$.
The construction of rotations is based on an analysis of stable assignments close to $x$ as exposed in the next lemma. It uses additional definitions.
Definition 3. Let $f\in F^{=}$. Define the reduced head for $f$ to be the set $\widetilde{H}_{f}=\widetilde{H}_{f}(x)$ consisting of non-saturated edges $e=fw$ in $H_{f}(x)$ such that $e$ belongs to $T_{w}(x)$. For each tie $\pi^{i}$ in $\Pi^{\rm end}_{f}(x)$, define the potential head for $f$ and $\pi^{i}$ to be the set $\widehat{H}_{f}^{i}=\widehat{H}_{f}^{i}(x)$ consisting of non-saturated edges $e=fw$ in $\pi^{i}$ such that $e$ belongs to $T_{w}(x)$, and no edge $e^{\prime}\in E_{f}$ with $e^{\prime}\geq_{f}e$ is incident to a deficit vertex in $W$. The best nonempty set among $\widetilde{H}_{f}(x)$ and $\widehat{H}_{f}^{i}$ ($\pi^{i}\in\Pi^{\rm end}_{f}(x)$) is denoted by $D_{f}=D_{f}(x)$ (ranging the sets according to the order in $\Pi_{f}$); if all these sets are empty, we put $D_{f}:=\emptyset$.
@@
Let $x^{\prime}\in{\cal S}$ be such that $x\succ x^{\prime}$ and $x^{\prime}$ is close to $x$, in the sense that the deviation ($\ell_{1}$-distance) $|x^{\prime}-x|=\sum(|x^{\prime}(e)-x(e)|\colon e\in E)$ is sufficiently small.
  • (i)
    Let $w\in W$ and $x^{\prime}_{w}\neq x_{w}$. Then $w$ is fully filled, and the following are valid:
    • (a)
      $H_{w}(x^{\prime})=H_{w}(x)$, and this is exactly the set of edges $e\in E_{w}$ such that $x^{\prime}(e)<x(e)$;
    • (b)
      no edge in $H_{w}(x)$ is incident to a deficit vertex in $F$.
  • (ii)
    Let $f\in F$ and $x^{\prime}_{f}\neq x_{f}$. Then $f$ is fully filled, and the following are valid:
    • (c)
      if $D_{f}(x)\neq\emptyset$, then $H_{f}(x^{\prime})=D_{f}(x)$, and this is exactly the set of edges $e\in E_{f}$ with $x^{\prime}(e)>x(e)$;
    • (d)
      no edge in $D_{f}$ is incident to a deficit vertex in $W$.
Proof Let $w\in W$ and $x^{\prime}_{w}\neq x_{w}$. Then $w\in W^{=}$ (by ((2.8))). The relation $x\succ_{F}x^{\prime}$ implies $x^{\prime}_{w}\succ_{w}x_{w}$ (by the polarity); therefore, the critical tie for $x^{\prime}_{w}$ is not worse than the one for $x_{w}$: $\pi^{\rm c}_{w}(x^{\prime})\geq\pi^{\rm c}_{w}(x)$ (by ((2.4))). Moreover, since $|x^{\prime}_{w}-x_{w}|=:\varepsilon>0$ is as small as wished, we can conclude that $\pi^{\rm c}_{w}(x^{\prime})=\pi^{\rm c}_{w}(x)$ and $H_{w}(x^{\prime})=H_{w}(x)$, yielding (a). (Here the inclusion $H_{w}(x^{\prime})\supseteq H_{w}(x)$ (valid by ((2.4))) turns into equality since we can choose $\varepsilon>0$ so that any $e\in\pi^{\rm c}_{w}(x)-H_{w}(x)$ and $a\in H_{w}(x)$ satisfy $x^{\prime}(e)-x(e)\leq\varepsilon$, $x(a)-x^{\prime}(a)\leq\varepsilon$, and $x(a)-x(e)>2\varepsilon$, implying $x^{\prime}(a)>x^{\prime}(e)$.) The assertion (b) follows from ((2.8)), taking into account that $x^{\prime}(e)\neq x(e)$ for all $e\in H_{w}(x)$.
Now consider $f\in F$ with $x^{\prime}_{f}\neq x_{f}$. Then $f\in F^{=}$ (by  ((2.8))) and $x_{f}\succ_{f}x^{\prime}_{f}$. Therefore, $\pi^{\rm c}_{f}(x)\geq\pi^{\rm c}_{f}(x^{\prime})$ (by ((2.4))). Suppose that $\pi^{\rm c}_{f}(x)=\pi^{\rm c}_{f}(x^{\prime})$. Then $H_{f}(x^{\prime})\subseteq H_{f}(x)$ (by ((2.4))). For any $e=fw\in H_{f}(x^{\prime})$, we have $x(e)<x^{\prime}(e)\leq b(e)$ (whence $e$ is not saturated by $x$), and $e$ belongs to $T_{w}(x^{\prime})=T_{w}(x)$ (taking into account that $H_{w}(x^{\prime})=H_{w}(x)$ and that $e$ cannot be in $H_{w}(x^{\prime})$). In case $H_{f}(x^{\prime})\neq H_{f}(x)$, any edge $e=fw$ in $H_{f}(x)-H_{f}(x^{\prime})$ belongs $T_{f}(x^{\prime})$, and therefore, $e$ must belong to $H_{w}(x^{\prime})=H_{w}(x)$ (since $x^{\prime}$ is stable). These reasonings imply that $H_{f}(x^{\prime})$ coincides with the reduced head $\widetilde{H}_{f}(x)$, whence $H_{f}(x^{\prime})=D_{f}(x)$ (for terminology and notation, see Definition 3).
Next suppose that $\pi^{\rm c}_{f}(x^{\prime})<\pi^{\rm c}_{f}(x)$. Let $\pi^{\rm c}_{f}(x^{\prime})$ be the tie $\pi^{i}$ in $\Pi_{f}$. Consider an arbitrary edge $e=fw$ in $E^{\rm beg}_{f}(x^{\prime})\cup(\pi^{i}-H_{f}(x^{\prime}))$. This edge belongs to the tail $T_{f}(x^{\prime})$, and since $x^{\prime}$ is stable, $e$ must belong to the head $H_{w}(x^{\prime})=H_{w}(x)$; in particular, the vertex $w$ is not deficit. It follows that $H_{f}(x^{\prime})$ is just the potential head $\widehat{H}^{i}_{f}(x)$ and that all sets $\widetilde{H}_{f}(x)$ and $\widehat{H}^{j}_{f}(x)$, where $\pi^{\rm c}_{f}(x)>\pi^{j}>\pi^{i}$, are empty. Therefore, $H_{f}(x^{\prime})=\widehat{H}^{i}_{f}(x)=D_{f}(x)$, yielding (c). In its turn, (d) follows from the fact that each edge $e=fw\in H_{f}(x^{\prime})$ satisfies $x(e)=0<x^{\prime}(e)$, whence $w$ is not deficit by ((2.8)). $\Box$
When this is not confusing, we will abbreviate notation $H_{v}(x),T_{v}(x),U_{v}(x)$ (where $x$ is fixed) to $H_{v},T_{v},U_{v}$. For $f\in F^{=}$, we call $D_{f}$ defined in (ii) of Lemma 4.1 the active head for $f$ and $x$.
Properties (b) and (d) in Lemma 4.1 can be generalized as follows:
@@
    let $(v_{0},e_{1},v_{1},\ldots,e_{k},v_{k})$ be a path such that: the vertices $v_{0},\ldots,v_{k-1}$ are fully filled while $v_{k}$ is deficit, and for $i=1,\ldots,k$, the edge $e_{k}$ belongs to $H_{v_{i-1}}$ (when $v_{i-1}\in W^{=}$) or $D_{v_{i-1}}$ (when $v_{i-1}\in F^{=}$); then $x^{\prime}(e_{i})=x(e_{i})$ holds for all $i$.
This is shown by induction on $k-i$. Since $v_{k}$ is deficit, we have $x^{\prime}(e_{k})=x(e_{k})$ and, moreover, $x^{\prime}(e)=x(e)$ for all edges $e\in E_{v_{k-1}}$. And for $i=0,\ldots,k-2$, the equality $x^{\prime}(e_{i})=x(e_{i})$ follows from $x^{\prime}(e_{i+1})=x(e_{i+1})$ and the balance condition $\sum(x^{\prime}(a)-x(a)\colon a\in E_{v_{i}})=0$, taking into account that for $a,b\in E_{v_{i}}$, the values $x^{\prime}(a)-x(a)$ and $x^{\prime}(b)-x(b)$ are equal if both $a,b$ belong to the corresponding head for $v_{i}$ and $x^{\prime}$, and have the same sign if both $a,b$ are not in this head.
This property motivates the following Cleaning procedure which is intended for extracting the vertices $v$ in $F^{=}\cup W^{=}$ where no changes of values on $E_{v}$ are possible.
@@
    We extract, step by step, singular vertices in $F^{=}\cup W^{=}$ as follows. In the beginning, we label as singular each vertex $v\in F^{=}\cup W^{=}$ whose head ($D_{v}$ or $H_{v}$) contains an edge incident to a deficit vertex. Also we label as singular each vertex $f\in F^{=}$ such that $D_{f}=\emptyset$ (since this should imply $x^{\prime}_{v}=x_{v}$). At a current step, if there appears a (new) vertex $v\in F^{=}\cup W^{=}$ whose head ($D_{v}$ or $H_{v}$) contains an edge incident to a vertex labeled as singular, then we label $v$ as singular as well.
The vertices in $F^{=}$ and $W^{=}$ not labeled as singular upon termination of this procedure are called regular, and the sets of these are denoted as $F^{+}$ and $W^{+}$, respectively. Note that
@@
    for each $f\in F^{+}$, we have $D_{f}\neq\emptyset$, and all edges $fw\in D_{f}$ are such that $w\in W^{+}$; similarly, for each $w\in W^{+}$, all edges $fw\in H_{w}$ are such that $f\in F^{+}$.
Also by Lemma 4.1(a),(c), the closeness of $x^{\prime}$ to $x$ (subject to $x\succ x^{\prime}$) corresponds to the property of homogeneity, namely:
@@
For $f\in F^{+}$, define $h_{f}:=|D_{f}|$, and for $w\in W^{+}$, define $h_{w}:=|H_{w}|$.
The reasonings above enable us to represent $\phi$ and $\psi$ defined in ((4.4)) via variables in the following system of equalities:
@@
(Here $\phi_{f}$ means a uniform increase of $x(e)$ at the edges $e\in D_{f}$, and $\psi_{w}$ a uniform decrease of $x(e)$ at the edges $e\in H_{w}$. The equalities in (4.5) describe the balance conditions for the vertices $f\in F^{+}$, and the ones in (4.6) describe similar conditions for the vertices $w\in W^{+}$, keeping the corresponding quota equalities.)
Also we should add to the above relations the following equality:
$\sum(h_{f}\phi_{f}\colon f\in F^{+})=\varepsilon$ (4.8)
for a sufficiently small number $\varepsilon>0$ (in order to avoid the equality $x^{\prime}=x$).
We fix such an $\varepsilon$, and our aim is to show that the set (polytope)
${\cal P}={\cal P}^{\varepsilon}(x):=\{(\phi,\psi)\}\subset{\mathbb{R}}^{F^{+}}\times{\mathbb{R}}^{W^{+}},$ (4.9)
where $(\phi,\psi)$ range the solutions to system (4.5)–(4.8), matches the set of augmentations $x^{\prime}-x$ for all $x^{\prime}\in{\cal S}$ such that $x^{\prime}\prec x$ and $x^{\prime}$ is close to $x$ (measured by the parameter $\varepsilon$).
More precisely, let us associate with $(\phi,\psi)\in{\cal P}$ the following vector $\Delta^{\phi,\psi}\in{\mathbb{R}}^{E}$:
@@
    for $e=fw\in E$, define $\Delta^{\phi,\psi}(e)$ to be equal to $\phi_{f}$ when $f\in F^{+}$ and $e\in D_{f}$, to be $-\psi_{j}$ when $w\in W^{+}$ and $e\in H_{w}$, and 0 otherwise.
Also define the set
$X^{\varepsilon}(x):=\{x+\Delta^{\phi,\psi}\colon(\phi,\psi)\in{\cal P}\}.$ (4.11)
@@
Let $x\in{\cal S}-\{x^{\rm max}\}$. Then for a sufficiently small $\varepsilon>0$,
(a) $X^{\varepsilon}(x)$ is exactly the set of stable assignments $x^{\prime}\in{\cal S}$ such that $x^{\prime}\prec x$ and the deviation $\sum(|x^{\prime}(e)-x(e)|\colon e\in E)$ is equal to $2\varepsilon$;
(b) if $y\in{\cal S}$ and $y\prec x$, then there exists $x^{\prime}\in X^{\varepsilon}(x)$ satisfying $y\preceq x^{\prime}\prec x$.
(Also we will explain that system (4.5)–(4.8) has no solution if and only if $x=x^{\rm max}$.)

### v2

@@ 3 Finding an initial stable assignment
In order to construct the stable matching $x^{\rm min}$ (which is the best for the part $F$ and the worst for the part $W$) in our mixed problem SMP, one can use a method due to Alkan and Gale applied in [1, Sec. 3.1] to show the existence of a stable assignment in a general case considered there. It iteratively constructs triples of functions $(b^{i},x^{i},y^{i})$ for $i=0,1,\ldots,i,\ldots$  .
@@
At the first stage of the iteration, $b^{i}$ is transformed into $x^{i}\in{\mathbb{R}}_{+}^{E}$, by putting $x^{i}_{f}:=C_{f}(b^{i}_{f})$ for each $f\in F$. (Then $x^{i}$ satisfies the quotas $q(f)$ for all “firms” $f\in F$, when dealing with our setting with quotas; however, $x^{i}$ can exceed quotas $q(w)$ for some “workers” $w\in W$.)
At the second stage, the obtained $x^{i}$ is transformed into $y^{i}\in{\mathbb{R}}_{+}^{E}$, by putting $y^{i}_{w}:=C_{w}(x^{i}_{w})$ for each $w\in W$. (This $y^{i}$ already satisfies the quotas $q(v)$ for all $v\in V$, but it need not be stable.)
The obtained functions $x^{i}$ and $y^{i}$ are then used to generate the upper bounds $b^{i+1}$ for the next iteration, by putting
@@
The process terminates if at some iteration, $i$-th say, we obtain the equality $y^{i}=x^{i}$ (which implies $b^{i+1}=b^{i}$, by (3.1)); then $x^{i}$ is just the required assignment (see the proposition below). In reality, it may happen that the sequence of iterations is infinite. Nevertheless, one can see (by arguing as in [1]) that the process always converges. Indeed, the construction shows that for each $i$,
@@
This result was proved in [1] for a general case (subject to axioms (A1)–(A3)); see Theorem 1 (for the stability) and Theorem 2 (for the optimality) in that paper.
Fortunately, for our mixed model the above method can be modified, by aggregating some series of consecutive iterations, so as to obtain an efficient algorithm to find the assignment $x^{\rm min}$; its running time is polynomial in $|E|$. We describe this modified method for SMP in the Appendix.
@@ 4 Rotations
Throughout this section, we consider a fixed stable assignment $x\in{\cal S}$ for our mixed problem SMP. Our aim is to characterize the set of stable assignments $x^{\prime}$ that are close to $x$ and satisfy $x\succ_{F}x^{\prime}$, namely, the set
$X^{\varepsilon}=X^{\varepsilon}(x):=\{x^{\prime}\in{\cal S}\colon x^{\prime}\prec_{F}x,\;|x^{\prime}-x|=\varepsilon\}$ (4.1)
for a sufficiently small real $\varepsilon>0$ (where $|x^{\prime}-x|$ denotes $\sum(|x^{\prime}(e)-x(e)|\colon e\in E)$). We will use a technique of rotations, generalizing a concept and constructions developed for earlier stability problems, starting with [12]. (In particular, it will be shown that $X^{\varepsilon}$ is a section of the polyhedral cone pointed at $x$ and generated by the rotations regarded as vectors in ${\mathbb{R}}^{E}$.) First of all we specify some notation needed in what follows.
$\bullet$ We denote by $V^{=}$ the set of fully filled vertices $v\in V$, which means that $|x_{v}|=q(v)$. The sets $V^{=}\cap F$ and $V^{=}\cap W$ are denoted by $F^{=}$ and $W^{=}$, respectively. (By ((2.7))(b), these sets do not depend on the choice of $x\in{\cal S}$. In view of ((2.8)), any transformation $x\rightsquigarrow x^{\prime}\in{\cal S}$ must involve only edges connecting $F^{=}$ and $W^{=}$.) The vertices in $V-V^{=}$ are called deficit.
@@
$\bullet$ An edge $e\in E$ is called saturated by $x$ if $x(e)=b(e)$, and unsaturated otherwise.
$\bullet$ When otherwise is not explicitly said, we write $\succeq$ for $\succeq_{F}$.
@@
    for each unsaturated edge $e=fw\in E$, at least one vertex $v$ among $f$ and $w$ is such that: $v$ is fully filled and $e$ belongs to $H_{v}(x)\cup E^{\rm end}_{v}(x)$.
The construction of rotations is based on an analysis of stable assignments close to $x$. It is convenient to assume that $\varepsilon$ is chosen so that:
  1. (4.3)
    $\varepsilon$ is less than any nonzero value among $x(e)$ and $b(e)-x(e)$ for $e\in E$, and less than $x(e)-x(e^{\prime})$ for edges $e\in H_{w}(x)$ and $e^{\prime}\in\pi^{\rm c}_{w}(x)-H_{w}(x)$ among all $v\in W^{=}$.
Then useful relations between $x$ and $x^{\prime}\in X^{\varepsilon}$ are described in the next two lemmas.
@@
Let $w\in W^{=}$ and $x^{\prime}_{w}\neq x_{w}$. Then $\pi^{\rm c}_{w}(x^{\prime})=\pi^{\rm c}_{w}(x)$,  $H_{w}(x^{\prime})=H_{w}(x)$, and $H_{w}(x)$ is exactly the set of edges $e\in E_{w}$ such that $x^{\prime}(e)<x(e)$.
Proof   By the polarity ((2.7))(b), $x^{\prime}\prec_{F}x$ implies $x^{\prime}_{w}\succ_{w}x_{w}$. Applying Lemma 2.2(ii) to $z:=x^{\prime}_{w}$ and $z^{\prime}:=x_{w}$, we have $\pi^{\rm c}_{w}(x^{\prime})\geq\pi^{\rm c}_{w}(x)$. If $\pi^{\rm c}_{w}(x^{\prime})>\pi^{\rm c}_{w}(x)$, then taking $e\in H_{w}(x)$, we obtain $\varepsilon=|x-x^{\prime}|\geq x(e)-x^{\prime}(e)=x(e)>\varepsilon$ (in view of ((4.3))); a contradiction. Therefore, $\pi^{\rm c}_{w}(x^{\prime})=\pi^{\rm c}_{w}(x)$ and $H_{w}(x^{\prime})\supseteq H_{w}(x)$. Moreover, $H_{w}(x^{\prime})=H_{w}(x)$ must take place (for otherwise, taking $e\in H_{w}(x)$ and $e^{\prime}\in H_{w}(x^{\prime})-H_{w}(x)$, we would have $\varepsilon\geq(x(e)-x^{\prime}(e))+(x^{\prime}(e^{\prime})-x(e^{\prime})=x(e)-x(e^{\prime})>\varepsilon$). Finally, for $e\in E_{w}-H_{w}(x)$, we have $x^{\prime}(e)\geq x(e)$ if $e\in T_{w}(x)$ (by Lemma 2.2(i)), and $x^{\prime}(e)=x(e)=0$ if $e\in E^{\rm end}_{w}(x)$, implying $H_{w}(x)=\{e\in E_{w}\colon x^{\prime}(e)<x(e)\}$. $\Box$
Definition 3. For $f\in F^{=}$, let $\widehat{\Pi}_{f}(x)$ denote the set of ties $\pi^{i}\in\Pi_{f}$ such that: (a) the set $D^{i}=D^{i}_{f}(x)$ of unsaturated edges $e=fw\in\pi^{i}$ such that $e\in T_{w}(x)$ is nonempty, and (b) there is no tie better or equal to $\pi^{i}$ that contains an unsaturated edge incident to a deficit vertex in $W$. When $\widehat{\Pi}_{f}(x)$ is nonempty, we denote the best (most preferred) tie $\pi^{i}$ in $\widehat{\Pi}_{f}(x)$ by $\widehat{\pi}_{f}(x)$, and denote $D^{i}$ by $D_{f}(x)$, called the potential head for $f$ and $x$. Otherwise we put $\widehat{\pi}_{f}(x):=\{\emptyset\}$ and $D_{f}(x):=\emptyset$. When $\widehat{\pi}_{f}(x)=\pi^{\rm c}_{f}(x)$, we also denote $D_{f}(x)$ as $\widetilde{H}_{f}(x)$ and call it the reduced head.
(Note that since $x$ is stable, each edge $e$ satisfying the condition in (a) (if exists) belongs to $H_{f}(x)\cup E^{\rm end}_{f}(x)$, cf. ((4.2)). Also if an edge $e=fw$ is unsaturated and $w$ is deficit, then $e$ cannot belong to $T_{f}(x)$. In case $\widehat{\pi}_{f}(x)=\pi^{\rm c}_{f}(x)$, we have $D_{f}(x)=\widetilde{H}_{f}(x)\subseteq H_{f}(x)$ and any edge $fw$ in $H_{f}(x)-\widetilde{H}_{f}(x)$ belongs to $T_{w}(x)$.)
Lemma 4.2
Let $f\in F^{=}$ and $x^{\prime}_{f}\neq x_{f}$. Then $\pi^{\rm c}_{f}(x^{\prime})\leq\pi^{\rm c}_{f}(x)$,  $H_{f}(x^{\prime})$ coincides with the set of edges $\{e\in E_{f}\colon x^{\prime}(e)>x(e)\}$, and if $\pi^{\rm c}_{f}(x^{\prime})=\pi^{\rm c}_{f}(x)$ then $H_{f}(x^{\prime})\subseteq H_{f}(x)$. Furthermore, $\pi^{\rm c}_{f}(x^{\prime})=\widehat{\pi}_{f}(x)$ and $H_{f}(x^{\prime})=D_{f}(x)$.
Proof  The inequality $\pi^{\rm c}_{f}(x^{\prime})\leq\pi^{\rm c}_{f}(x)$ and the inclusion $H_{f}(x^{\prime})\subseteq H_{f}(x)$ in case $\pi^{\rm c}_{f}(x^{\prime})=\pi^{\rm c}_{f}(x)$ follow from Lemma 2.2(ii) with $z:=x_{f}$ and $z^{\prime}:=x^{\prime}_{f}$. In order to see the equality $H_{f}(x^{\prime})=\{e\in E_{f}\colon x^{\prime}(e)>x(e)\}$, observe that for an edge $e=fw$, if $x^{\prime}(e)>x(e)$, then $x^{\prime}(e)\leq x(e)+\varepsilon<b(e)$ (in view of ((4.3))) and there holds $e\in T_{w}(x)=T_{w}(x^{\prime})$ (by Lemma 4.1). Therefore, $e\notin T_{f}(x^{\prime})$ (since $x^{\prime}(e)<b(e)$ and $x^{\prime}$ is stable, cp. ((4.2))). This implies $e\in H_{f}(x^{\prime})$ (taking into account that $e\in E^{\rm end}_{f}(x^{\prime})$ is impossible since $x^{\prime}(e)>0$).
Finally, by the stability of $x^{\prime}$, there is no edge $e=fw$ in $E^{\rm beg}_{f}(x^{\prime})\cup(\pi^{\rm c}_{f}-H_{f}(x^{\prime}))$ such that $x^{\prime}(e)<b(e)$ and $e\in T_{w}(x^{\prime})$ ($=T_{w}(x)$) (in particular, $w$ cannot be deficit). Also $H_{f}(x^{\prime})$ is nonempty and each edge $e$ in it is unsaturated by $x$ (since $x(e)<x^{\prime}(e)$). Using these facts, one can conclude that $\pi^{\rm c}_{f}(x^{\prime})=\widehat{\pi}_{f}(x)$ and $H_{f}(x^{\prime})=D_{f}(x)$. $\Box$
Lemmas 4.1 and 4.2 together with property ((2.8)) imply that
@@
    for $x^{\prime}\in X^{\varepsilon}$, if $w\in W^{=}$ and $x^{\prime}_{w}\neq x_{w}$, then no edge in $H_{w}(x)$ is incident to a deficit vertex in $F$; and if $f\in F^{=}$ and $x^{\prime}_{f}\neq x_{f}$, then $D_{f}(x)\neq\emptyset$.
This can be generalized as follows.
Lemma 4.3
Let $(v_{0},e_{1},v_{1},\ldots,e_{k},v_{k})$ be a path in $G$ such that: the vertices $v_{0},\ldots,v_{k-1}$ are fully filled, and for $i=1,\ldots,k$, the edge $e_{i}$ connecting $v_{i-1}$ and $v_{i}$ belongs to $H_{v_{i-1}}(x)$ (when $v_{i-1}\in W^{=}$) or $D_{v_{i-1}}(x)$ (when $v_{i-1}\in F^{=}$). Suppose that either $v_{k}\in F-F^{=}$, or $v_{k}\in W$ and $D_{v_{k-1}}(x)=\emptyset$. Then any $x^{\prime}\in X^{\varepsilon}(x)$ and $i=1,\ldots,k$ satisfy $x^{\prime}(e_{i})=x(e_{i})$.
Proof  This is shown by induction on $k-i$. If $v_{k}\in F$ is deficit, we have $x^{\prime}(e_{k})=x(e_{k})$ (by ((2.8))), implying $x^{\prime}(e)=x(e)$ for all edges $e\in E_{v_{k-1}}$ (since $e_{k}\in H_{w}(x^{\prime})=H_{w}(x)$). And if $v_{k}\in W$ and $D_{v_{k-1}}(x)=\emptyset$, then $x^{\prime}_{v_{k-1}}=x_{v_{k-1}}$. See ((4.4)). Now for $i=0,\ldots,k-2$, the equality $x^{\prime}(e_{i})=x(e_{i})$ follows from $x^{\prime}(e_{i+1})=x(e_{i+1})$ and the balance condition $\sum(x^{\prime}(a)-x(a)\colon a\in E_{v_{i}})=0$, taking into account that for $a,b\in E_{v_{i}}$, the values $x^{\prime}(a)-x(a)$ and $x^{\prime}(b)-x(b)$ are equal if both $a,b$ belong to the corresponding head ($H_{v_{i}}(x)$ or $D_{v_{i}}(x)$), and have the same sign if both $a,b$ are not in this head. $\Box$
In what follows, since $x$ is fixed, we may abbreviate notation $H_{v}(x),T_{v}(x),D_{v}(x)$ to $H_{v},T_{v},D_{v}$.
Lemma 4.3 motivates the following Cleaning procedure which is intended to extract a set $V^{0}$ of vertices $v\in F^{=}\cup W^{=}$ where no changes of $x$ on $E_{v}$ are possible.
@@
    We grow $V^{0}$, step by step, as follows. Initially, relying on ((4.4)), we insert in $V^{0}$ each vertex $w\in W^{=}$ whose head $H_{v}$ contains an edge incident to a deficit vertex, and each vertex $f\in F^{=}$ whose potential head $D_{f}$ is empty. At a current step, if there appears a (new) vertex $v\in F^{=}\cup W^{=}$ whose head ($D_{v}$ or $H_{v}$) has an edge incident to a vertex already contained in $V^{0}$, then we insert $v$ in $V^{0}$ as well.
The fully filled vertices $v\in V^{=}$ occurring in $V^{0}$ upon termination of this procedure are called singular, and the remaining fully filled vertices are called regular. We denote $V^{+}:=V^{=}-V^{0}$,  $F^{+}:=V^{+}\cap F$ and $W^{+}:=V^{+}\cap W$. Note that
@@
    for each $f\in F^{+}$, we have $D_{f}\neq\emptyset$, and all edges $fw\in D_{f}$ satisfy $w\in W^{+}$; similarly, for each $w\in W^{+}$, all edges $fw\in H_{w}$ satisfy $f\in F^{+}$.
By reasonings above, all transformations $x\rightsquigarrow x^{\prime}\in X^{\varepsilon}(x)$ can be performed only within the sets $H_{w}$ for regular $w\in W^{+}$ and $D_{f}$ for regular $f\in F^{+}$; we call the edges in these sets active. (The heads $H_{w}$ and $D_{f}$ may be called active as well.)
(As we have seen, the construction of active sets for $F^{+}$ looks more involved compared with those for $W^{+}$. The reason of this asymmetry is caused by their different roles in a small transformation of $x$; namely, to provide more preferred assignments in the part $W$, and less preferred in the part $F$, while preserving stability, we should slightly decrease $x$ within full current heads for $W$, but increase it either in reduced heads for $F$, or even in ties later than the critical ones where all values are zeros.)
Next we will associate with the regular set $V^{+}$ two important structures: a graph generated by active edges, and a system of linear equations. We first prefer to describe the second structure.
By Lemmas 4.1 and 4.2, the closeness of $x^{\prime}$ to $x$ subject to $x\succ_{F}x^{\prime}$ corresponds to the property of homogeneity, namely:
@@
These values $\phi_{f}$ and $\psi_{w}$ serve as variables in the following system where we use as coefficients the scalars $h_{f}:=|D_{f}|$ and $h_{w}:=|H_{w}|$:
@@
Here $\phi_{f}$ means a uniform increase of $x(e)$ at the edges $e\in D_{f}$, and $\psi_{w}$ a uniform decrease of $x(e)$ at the edges $e\in H_{w}$. (Recall that for $x^{\prime}\in X^{\varepsilon}$ with $\varepsilon$ small, $H_{f}(x^{\prime})=D_{f}(x)$ for all $f\in F^{+}$, and $H_{w}(x^{\prime})=H_{w}(x)$ for all $w\in W^{+}$.) The equalities in (4.7) describe the balance conditions for the vertices $f\in F^{+}$ (saying that the total increase of $x$ on $D_{f}$ is compensated by the total decrease on $E_{f}-D_{f}$, keeping the corresponding quota fillings), and the ones in (4.8) describe similar balance conditions for the vertices $w\in W^{+}$. And (4.9) expresses the condition $|x^{\prime}-x)=\varepsilon$.
Our aim is to show that the set (polytope)
${\cal P}^{\varepsilon}={\cal P}^{\varepsilon}(x):=\{(\phi,\psi)\}\subset{\mathbb{R}}^{F^{+}}\times{\mathbb{R}}^{W^{+}},$ (4.11)
where $(\phi,\psi)$ runs through the solutions to system (4.7)–(4.10), matches the set of augmentations $x^{\prime}-x$ for all $x^{\prime}\in X^{\varepsilon}(x)$.
More precisely, let us associate with $(\phi,\psi)\in{\cal P}^{\varepsilon}$ the following vector $\Delta^{\phi,\psi}\in{\mathbb{R}}^{E}$:
@@
    for $e=fw\in E$, define $\Delta^{\phi,\psi}(e):=\phi_{f}$ if $f\in F^{+}$ and $e\in D_{f}$,  $-\psi_{j}$ if $w\in W^{+}$ and $e\in H_{w}$,  and 0 otherwise.
@@
Let $x\in{\cal S}-\{x^{\rm max}\}$. Then:
(a) for a sufficiently small real $\varepsilon>0$ (in particular, satisfying ((4.3))), the set $X^{\varepsilon}(x)$ defined in (4.1) coincides with the polytope $\{x+\Delta^{\phi,\psi}\colon(\phi,\psi)\in{\cal P}^{\varepsilon}\}$;
(b) if $y\in{\cal S}$ and $y\prec x$, then for a sufficiently small $\varepsilon>0$, there exists $x^{\prime}\in X^{\varepsilon}(x)$ satisfying $y\preceq x^{\prime}\prec x$.
(Also we shall see that system (4.7)–(4.10) has no solution if and only if $x=x^{\rm max}$.)